OpenStudy (anonymous):

a microscope slide shows 45 red blood cells out of 75 blood cells. how many red blood cells would be expected in a sample of the same blood that has 450 cells

OpenStudy (anonymous):

Come on atleast tell her how ha Find the ratio: 45/75 = 0.6......Then multiply 450*0.6 = 270 red blood cells

OpenStudy (luigi0210):

I think you could use: \[\LARGE \frac{45}{75}=\frac{x}{450}\] Or yea, her way seems easier.

OpenStudy (anonymous):

thnx

OpenStudy (luigi0210):

btw, \[\LARGE \color{green}{WELCOME~~TO~~OPENSTUDY}\]

OpenStudy (anonymous):

thnx